# A Tale of Tails: Model Collapse as a Change of Scaling Laws

### v1

@@ 2.3 Collapse Over Multiple Generations of AI Data
$\begin{split}E^{(n)}_{test}&\asymp T^{-c}+\underbrace{T_{0}^{-c}+\ldots+T_{0}^{-c}}_{n\text{ times}}\\ &=T^{-c}+C_{n}T_{0}^{-c}=T^{-c}\left(n(T/T_{0})^{c}+1\right),\end{split}$ (12)
where $c:=1-1/\beta$. We deduce the following result.
@@
$E^{(n)}_{test}\asymp C_{n}T^{-c}\asymp{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}n}T^{-c}.$ (13)
@@ 3 Mitigating Model Collapse via Data Mixing
Here we explore the possibility of alleviating model collapse via the acquisition of even a tiny amount of data from the true data distribution, to complement AI polluted data. We study two phenomena: (1) In the case of mixing $\pi$-fraction of the original data with a $(1-\pi)$ fraction of AI-generated data we exhibit a startling ”grokking” phenomenon where test loss plateaus with increasing training data to finally decrease again according to the scaling law of the original model, and (2) in the scenario where we would like to compensate for missing ”tail”, we acquire some data from the tail of the original distribution to show that this needs to be done with caution: getting data from ”too deep” in the tail is worthless while data closer to the precise ”missing” tail can be beneficial. All proofs can be found in Appendix C.
@@ 3.1 Acquiring Missing Tail
To counter the effect of tail cutting and the resulting plateau in scaling, we might resort to adding curated data that would emphasize the tail. The following Theorem 3.1 studies this effect; it shows, in particular that if we ”overshoot” and only curate tail that is too deep, our efforts will be worthless. Rather, there is a fine line around the chopped tail $k$ (within a factor of $(1+o(1))$ of $k$), where we need to place our data curation efforts to achieve the desired effect, a return of the scaling beyond the plateau.
@@
(B) If $C>1$, then $E_{test}\asymp T^{-c}+k^{-\alpha}$ (which recovers the result of Theorem 2.1), and so ”buying” the $N$th tail of the real data distribution is worthless.
@@ 3.2 A Grokking Phenomenon
Here we show how even small amounts of original data can mitigate the above ”scaling law collapse” by introducing a grokking phenomenon where test error plateaus and eventually continues to decline.
@@ 4.3 Model Collapse in Probabilistic Setting
Grokking for Mixtures. Technically speaking, this grokking phenomenon only holds for models with deterministic ground truth labels, like the Hutter LLM and the limited capacity associative memory model. For the probabilistic setting of bigrams (or text LLMs) the theorem cannot hold in its pure form, because if we train on a mixture of two distributions (clean and synthetic) but test only on the clean distribution, the distance between these two distributions will always be a lower bound on the test error. However, we can see that remnants of a ”smoothed” grokking-law persist in the form of an S-shaped scaling (see Figure 19 in Appendix B).
@@ 5 Capacity-Limited Memory Models: A Triplet Scaling Law
We now turn to a finite-memory extension of the Hutter LLM, which allows to model capacity. We thus study the model collapse phenomenon in the context of the following simple associative memory model studied in (Cabannes et al., 2023)
$\begin{split}f_{T}(i)&:=\arg\max_{y}H_{T}(i,y),\text{ where}\\ H_{T}(i,y)&:=e_{i}^{\top}M_{T}u_{y},\\ M_{T}&:=\sum_{i}q_{T}(i)e_{i}u_{f_{\star}(i)}^{\top}\in\mathbb{R}^{d\times d}.\end{split}$ (18)
@@
$E_{test}\asymp T^{-(1-1/\beta)}+d^{-c_{q}}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}k^{-(\beta-1)}}.$ (21)
@@ 6 Experiments
In our theoretical analysis, motivated by empirical observations (see Figure 2) or by the effect of finite data sampling bias on heavy-tailed data, we have assumed that generated data follows patterns of either a tail cutoff or tail narrowing. In our subsequent experiments, we depart from theoretical assumptions on tail-cutting/narrowing to allow the widely deployed top-p selection or temperature scaling mechanisms to give possibly intermingled effects on the generated data distribution.
@@
Figure 6: Hutter++ on Bigram with limited data and top-p. The initial model is trained on $T_{0}=100,000$ samples. It generates $T$ samples for Gen 1. Starting from Gen 2 models are trained on data generated by the most powerful model from the previous generation. Top-p-0.95 cutting and $\beta=3/2$.
@@
Autoregressive Bigram Models with Perplexity Loss. We move one step further towards ”real” LLMs to investigate autoregressive bigram models. The dataset now comprises sequentially generated integers, adhering to Equation (16), with the model trained on all tokens. We use the averaged perplexity score of the test set as the test error metric. Our study encompasses a range of effects—such as top-p inference, temperature scaling, limited real data, and training on progressively larger AI datasets. Consistent with the findings in Section 4, we observe the same patterns of scaling loss and progressive model collapse across generations. Relevant figures are provided in Appendix F.
@@
Transformers Learning the GCD. Our first illustration of our theory ”in the wild” is for sequence-to-sequence transformer models for an arithmetic task: predicting the greatest common divisor (GCD) of two integers, encoded as sequences of digits in some base $B$ following Charton (2023). This setup is a perfect intermediate step between our toy models and large scale LLMs; it uses the transformer architecture and training algorithms on sizeable models, while the underlying data has a deterministic nature. Over the course of training the model progressively learns new GCDs and with them also their products with already learned GCDs. We can thus view each such learned group, usually learned in ”bursts”, as a new skill. For the purpose of this experiment, we use this model after $300M$ samples as the generator of AI-data. In Figure 4 we validate the predicted scaling law for a single generation and observe ‘un-learning’ of skills when training exclusively with generated data, as well as a grokking effect when training with mixtures. See Appendix G for a full description and more details and Figures.
@@
Experiments on LLMs. We finetune Llama2 with LoRA, generating synthetic AI data for the next finetuning iteration. Inspired by the setup in Shumailov et al. (2023), we use Wikidata-103, partitioned into approximately $2.2$ million sequences of 128 tokens. AI data is generated through prompt completion, using the first 96 tokens from the original sequences as prompts. The model is trained only on the last 32 tokens to preclude information leakage, i.e. the model being trained on the ground truth of the same 32 tokens. The evaluations are conducted exclusively on the same 32 tokens. We use top-p 0.9 and temperature 0.9 across all generation. The results, depicted in Figure 4 (left), illustrate a scaling law decay over several generations. The first generated dataset still contain useful but limited information and the utility of the second generation’s data markedly diminishes. These phenomena corroborate the anticipated loss of scaling law and model collapse, further indicating that model collapse is even more pronounced here, highlighting the challenges in training next generation LLMs. More details and results in Appendix H.
@@
Figure 7: Mixing Llama Generated Data with Original Data Based on Figure 4 left, we further mix the generated data with original data, with ratio 98 to 2. Adding original data significantly mitigates the model collapse. Look how the mixing curve validates our predicted curve of the grokking phenomenon as in Figure 3
@@ 7 Conclusion
In the advent of the ”synthetic data age”, our work signals the end of current neural scaling laws and opens the door to a puzzling array of new phenomena in a world where the training corpora are enriched with AI generated data. We demonstrate that scaling laws cease to persist; test error tapers off due to altered, less heavy tailed, data distributions. As noted already in prior work, in a fully synthetic data world learning will stop and models will degenerate - their scaling will halt and revert completely. Yet, new opportunities arise from careful mixture and data curation, as we have shown, with interesting effects at the interplay of clean and synthesized data. We must recognize new learning plateaus and, for instance, adjust to changed learning curves from blending clean and synthetic data to unintended early stopping. A notable feature of our work is that our theory is effective - we observe the predicted phenomena for relevant large models in two different settings.
Taken together, our contributions call for a more responsible, or ”collapse-aware”, proliferation of synthesized data. Scale is not all you need: more work on effective watermarking for synthetic data is needed, to make it more distinguishable from the original, human-annotated data. Thus, clean / real data will become an even more valuable resource in the future, as we are ushering in the ”beyond scaling” era.
8 Acknowledgements
@@ Appendix A Prior Work
The phenomenon appeared in the recent literature in the context of language and image generation. Several recent works demonstrate facets of this phenomenon empirically in various settings (Hataya et al., 2023; Martínez et al., 2023a, b; Bohacek & Farid, 2023; Briesch et al., 2023; Guo et al., 2023; Fan et al., 2023). Only few recent works also provide some accompanying theoretical analysis (Shumailov et al., 2023; Alemohammad et al., 2023; Bertrand et al., 2023) which we outline now.
@@
(Shumailov et al., 2023) define model collapse and attribute it to two mechanisms: finite sampling when training a model (leading to cut off of low-probability data) and function approximation errors (the model is not sufficiently expressive to model the true distribution). They observe (and, for a single Gaussian, prove) that upon iteratively resampling finite ”training data” the generated distribution becomes more and more peaked. Other models studied empirically are mixtures of (two) Gaussians and VAEs on MNIST. To study language models, (Shumailov et al., 2023) iteratively fine tune Meta’s OPT-125M model on wikidata2. For generation of new text they use a 5-way beam search, which, by its nature, (approximatively) generates only low-perplexity data.
@@
(Alemohammad et al., 2023) conduct an empirical and analytical analysis on generative image models of what they term the ”self-consuming” or ”autophaguous” loop. They conclude that without enough fresh real data at each generation, future models necessarily will have their precision or recall decrease. They model the influence of each new AI-generation via a generic sampling bias $0\leq\lambda\leq 1$. In the case of image generation this refers to feature parameters at generation that favor quality over diversity (suitably quantified). More precisely, $\lambda=1$ corresponds to unbiased sampling and $\lambda=0$ corresponds to sampling from the modes of the generative distribution. $\lambda$ models biased sampling methods commonly used in generative modeling practice, such as truncation in BigGAN and StyleGAN or guidance in diffusion models. In the case of Gaussian distributions, $\lambda$ is the shrinking factor of the variance of the next generation. Their empirical work studies GANs and denoising diffusion probabilistic models for image generation on FFHQ and MNIST and single Gaussians for both theoretical and empirical observations. As in (Shumailov et al., 2023) they observe (and prove for the case of a single Gaussian) that estimation error alone leads to vanishing variance with number of iterations. (Alemohammad et al., 2023) also empirically observe an initial boost in performance in a regime where modest amounts of synthetic data are mixed with the original data before larger amounts of synthetic data lead to ultimate degradation. This might mimick larger-scale results that demonstrate how synthetic data mixed with true data improves performance in some scenarios (see Benefits of synthesized data below). Indeed, in its simplest form, data augmentation (rotations, cropping etc. ), a widespread highly beneficial practice in ML training, can be viewed as the simplest form of data generation.
@@
Let us mention the study of Bertrand et al. (2023) in the context of image generation, which sheds light on the critical role of data composition in the stability and effectiveness in generative models. They explore scenarios involving a mix of clean data, representative of the true distribution, and synthesized data from previous iterations of the generator. Their analysis reveals that if the data mix consists exclusively of synthesized data, the generative process is likely to degenerate over time, leading to what they describe as a ’clueless generator’. Thus, the generator collapses: it progressively loses its ability to capture the essence of the data distribution it was intended to model. Conversely, they found that when the proportion of clean data in the mix is sufficiently high, the generator, under certain technical conditions, retains the capability to learn and accurately reflect the true data distribution. This work sheds light on the critical role of data composition in the stability and effectiveness of generative models.
@@
Several empirical studies confirm the deleterious effect of training on self-generated data: In the context of image generation, (Martínez et al., 2023a, b) report degradation of models trained on AI-generated data. Specifically, they use a Denoising Diffusion Implicit Model and a few (relatively small) datasets (e.g. Orchids, MNIST) to demonstrate visual degradation when training in successive generations of AI-generated data. (Hataya et al., 2023) ”conclude that generated images negatively affect downstream performance, while the significance depends on tasks and the amount of generated images”, (Bohacek & Farid, 2023) reports that the popular StableDiffusion model collapses when iteratively retrained on self-generated faces, even with as little as $3\%$ synthetic data mixed into the original training set. For text, (Briesch et al., 2023) use nanoGPT [^5] https://github.com/karpathy/nanoGPT on a curated 10K logical-expression dataset to demonstrate the iterative collapse of self-consuming loops - the model and dataset are sufficiently small to allow training from scratch. (Guo et al., 2023) observe a decline in linguistic diversity metrics across iteratively fine-tuned LLMs.
@@
To our knowledge, rigorous theory (or even empirical demonstrations) on mitigation strategies against model collapse are yet to come, with one notable exception in (Bertrand et al., 2023) (see below). Several works discuss the need for detection of AI-generated images or text (to avoid retraining on them), for example motivating research into watermarking strategies. (Bertrand et al., 2023) analyze iterative retraining on a mixture of synthesized and original data under several technical assumptions and find that there are fixed points governing the stability of iterative retraining.
@@
There is a range of results showing benefits of AI-synthesized data in training better models, though mostly these results pertain to image data, specifically in the context of diffusion models (Azizi et al., 2023; He et al., 2023; Shipard et al., 2023; Bansal & Grover, 2023; Lin et al., 2023), though not only (see Dai et al. (2023); Xu et al. (2023); Huang et al. (2022); Wang et al. (2023) for chat-related examples). One might argue that they either throw model-collapse caution to the winds or, possibly, settle in the protected corner where mild amounts of synthetic data (or larger amounts of ”mildly synthetic” data, like in the case of data augmentation) helps. In particular, often benefits of synthetic data are observed when the synthetic data is generated by a model trained for a different use case than the downstream task (like images synthesized from diffusion models helping classification models) or generated by a stronger model (He et al., 2023; Shipard et al., 2023; Bansal & Grover, 2023; Lin et al., 2023). However, other works critically analyze the purported benefit of generated data. (Burg et al., 2023) find that while synthesized data from a diffusion model helps improving downstream tasks, such as classification, using the pre-training data of the diffusion model alone gives even stronger performance (which we can interpret as evidence of mild first-generation model collapse). All in all it is fair to say that the impact of data augmentation using generative models is still not fully understood.
@@
This study was extended to demonstrate many more power law relations in various scenarios (vision transformer, video modeling, multimodal models, and mathematical problem solving) (Henighan et al., 2021). In the machine translation (MT) setting, (Gordon et al., 2021) quantify scaling laws for standard benchmarks like BLEU and explain them via cross-entropy power-law scaling, thus positing a first universality of scaling laws across metrics. (Hernandez et al., 2021) demonstrate similar empirical power-law scaling for transfer learning and (Aghajanyan et al., 2023) provide a vast experimental body of evidence for scaling laws in mixed-modal language models.
However, a few results have nuanced the view of scaling as a panacea to improved loss. For instance, (McKenzie et al., 2023) present evidence for ”inverse sclaing” where flaws in the training objective or the data lead to U-shaped scaling.
@@
From a theoretical angle, scaling laws have been shown analytically even before the emergence of large foundation models. For instance, Caponnetto & de Vito (2007) characterize the power-law generalization error of regularized least-squares kernel algorithms. The role of optimization can also be taken into account in this setting ((Nitanda & Suzuki, 2021)). In the nonparametric literature, for example (Schmidt-Hieber, 2017) and (Suzuki, 2019) derived the test error scaling of deep neural network in fitting certain target functions and (Bordelon et al., 2020) analyze spectral dependence.
More recently, scaling laws have been shown for kernel models under the Gaussian design, e.g. in (Spigler et al., 2020; Cui et al., 2021, 2022) for regression and (Cui et al., 2023) for classification. (Maloney et al., 2022) study scaling laws for the random feature model in the context of regression. In the context of memorization for heavy-tailed data scaling laws have been shown in the infinite-memory setting (Hutter, 2021), for ”quantized” skills (Michaud et al., 2023) and for certain random data-generation processes (Debowski, 2023). When taking model capacity and optimization into account, Cabannes et al. (2023) recently proved scaling laws in constraint-capacity associative memories.
@@
Scaling laws give us an insight on bang-for-the-buck style trade-off for model training. However, cross-entropy loss is not a goal in and of itself: we want to train models that are endowed with a larger and larger skill set as we scale them up. For instance, (Gordon et al., 2021) provide intuition and empirics for the scaling of BLEU score for MT with cross-entropy loss as

### v2

@@ 2.3 Collapse Over Multiple Generations of AI Data
$\begin{split}E^{(n)}_{test}&\asymp T^{-c}+\underbrace{T_{0}^{-c}+\ldots+T_{0}^{-c}}_{n\text{ times}}\\ &=T^{-c}+nT_{0}^{-c}=T^{-c}\left(n(T/T_{0})^{c}+1\right),\end{split}$ (12)
where $c:=1-1/\beta$. Only in the context of model collapse across multiple generations, $T$ is the size of the AI data in the final step and $T_{0}$ is the data a model trained in all preceding steps. We deduce the following result.
@@
$E^{(n)}_{test}\asymp C_{n}T^{-c}\asymp{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}n}T^{-c}.$ (13)
@@ 3 Mitigating Model Collapse via Data Mixing
Here we explore the possibility of alleviating model collapse via the acquisition of even a tiny amount of data from the true data distribution, to complement AI polluted data. We study two phenomena: (1) In the case of mixing $\pi$-fraction of the original data with a $(1-\pi)$ fraction of AI-generated data we exhibit a startling “grokking” phenomenon where test loss plateaus with increasing training data to finally decrease again according to the scaling law of the original model, and (2) in the scenario where we would like to compensate for missing “tail”, we acquire some data from the tail of the original distribution to show that this needs to be done with caution: getting data from “too deep” in the tail is worthless while data closer to the precise “missing” tail can be beneficial. All proofs can be found in Appendix C.
@@ 3.1 Acquiring Missing Tail
To counter the effect of tail cutting and the resulting plateau in scaling, we might resort to adding curated data that would emphasize the tail. The following Theorem 3.1 studies this effect; it shows, in particular that if we “overshoot” and only curate tail that is too deep, our efforts will be worthless. Rather, there is a fine line around the chopped tail $k$ (within a factor of $(1+o(1))$ of $k$), where we need to place our data curation efforts to get the scaling back.
@@
(B) If $C>1$, then $E_{test}\asymp T^{-c}+k^{-\alpha}$ (which recovers the result of Theorem 2.1), and so “buying” the $N$th tail of the real data distribution is worthless.
@@ 3.2 A Grokking Phenomenon
Here we show how even small amounts of original data can mitigate the above “scaling law collapse” by introducing a grokking phenomenon where test error plateaus and eventually continues to decline.
@@ 4 A Tailed Bigram Model
Remark.
The Hutter model abstracts Large Language Models (LLMs) into encapsulated knowledge or skills Michaud et al. (2023), denoted as $i$. This variable, $i$, can be interpreted as a composite description comprising multiple components, where the Hutter model highlights the tendency of AI-generated data to exhibit a less pronounced heavy-tailed distribution. In scenarios where AI-generated data encompasses both accurate and erroneous information—such as in predictions of the greatest common divisor—a bias towards incorrect responses is observed. The Hutter++ model extends this framework by incorporating this specific bias. Specifically, we now curate the model to generate AI data with the consistent probability distribution $p(i))$. Errors are represented within the conditional distribution $p(j|i)$, which, in turn, affects the model’s performance in later training stages.
@@ 4.3 Model Collapse in Probabilistic Setting
Grokking for Mixtures. Technically speaking, this grokking phenomenon only holds for models with deterministic ground truth labels, like the Hutter LLM and the limited capacity associative memory model. For the probabilistic setting of bigrams (or text LLMs) the theorem cannot hold in its pure form, because if we train on a mixture of two distributions (clean and synthetic) but test only on the clean distribution, the distance between these two distributions will always be a lower bound on the test error. However, we can see that remnants of a “smoothed” grokking-law persist in the form of an S-shaped scaling (see Figure 19).
@@ 5 Capacity-Limited Memory Models: A Triplet Scaling Law
We now turn to a finite-memory extension of the Hutter LLM, which allows to model capacity. We look into a simple associative memory model (Cabannes et al., 2023):
$\begin{split}f_{T}(i)&:=\arg\max_{y}H_{T}(i,y),\text{ where}\\ H_{T}(i,y)&:=e_{i}^{\top}M_{T}u_{y},\\ M_{T}&:=\sum_{i}q_{T}(i)e_{i}u_{f_{\star}(i)}^{\top}\in\mathbb{R}^{d\times d}.\end{split}\vspace{-0.3cm}$ (18)
@@
$E_{test}\asymp T^{-(1-1/\beta)}+d^{-c_{q}}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}k^{-(\beta-1)}}.$ (21)
@@ 6 Experiments
In our theoretical analysis, motivated by empirical observations (see Figure 2) or by the effect of finite data sampling bias on heavy-tailed data, we have assumed that generated data follows patterns of either a tail cutoff or tail narrowing. In our subsequent experiments, we depart from theoretical assumptions on tail-cutting/narrowing to allow the widely deployed top-$p^{inf}$ selection or temperature scaling mechanisms to give possibly intermingled effects on the generated data distribution.
@@
Figure 6: Hutter++ on Bigram with limited data and top-$p^{inf}$. The initial model is trained on $T_{0}=100,000$ samples. It generates $T$ samples for Gen 1. Starting from Gen 2 models are trained on data generated by the most powerful model from the previous generation. Top-$p^{inf}=$0.95 cutting and $\beta=3/2$.
@@
Autoregressive Bigram Models with Perplexity Loss. We move one step further towards “real” LLMs to investigate autoregressive bigram models. The dataset now comprises sequentially generated integers, adhering to Equation (16), with the model trained on all tokens. We use the averaged perplexity score of the test set as the test error metric. Our study encompasses a range of effects—such as top-$p^{inf}$ inference, temperature scaling, limited real data, and training on progressively larger AI datasets. Consistent with the findings in Section 4, we observe the same patterns of scaling loss and progressive model collapse across generations. Relevant figures are provided in Appendix F.
@@
Transformers Learning the GCD. Our first illustration of our theory “in the wild” is for sequence-to-sequence transformer models for an arithmetic task: predicting the greatest common divisor (GCD) of two integers, encoded as sequences of digits in some base $B$, following Charton (2023). This setup is a perfect intermediate step between our toy models and large scale LLMs; it uses the transformer architecture and training algorithms on sizeable models, while the underlying data has a deterministic nature. Over the course of training the model progressively learns new GCDs and with them also their products with already learned GCDs. We can thus view each such learned group, usually learned in “bursts”, as a new skill. For the purpose of this experiment, we use this model after $300M$ samples as the generator of AI-data. In Figure 4 we validate the predicted scaling law for a single generation and observe ‘un-learning’ of skills when training exclusively with generated data, as well as a grokking effect when training on mixtures. See Appendix G for details and more figures.
@@
Experiments on LLMs. We finetune Llama2 with LoRA, generating synthetic AI data for the next finetuning iteration. Inspired by the setup in Shumailov et al. (2023), we use Wikidata-103, partitioned into approximately $2.2$ million sequences of 128 tokens. AI data is generated through prompt completion, using the first 96 tokens from the original sequences as prompts. The model is trained only on the last 32 tokens to preclude information leakage. The evaluations are conducted exclusively on the same 32 tokens. We use top-$p^{inf}=$0.9 and temperature $\tau=$0.9 across all generations. The results, depicted in Figure 4 (left), illustrate a scaling law decay over several generations. The first generated dataset still contains useful but limited information; the utility of the second generation’s data markedly diminishes. These phenomena corroborate the anticipated loss of scaling law and model collapse, further indicating that model collapse is even more pronounced here, highlighting the challenges in training next generation LLMs. More details and results in Appendix H.
@@
Figure 7: Mixing Llama Generated Data with Original Data Based on Figure 4 left, we mix generated data with original data, with ratio 98 to 2. Note how the mixing curve validates our predicted curve of the grokking phenomenon as in Figure 3
@@ 7 Conclusion
In the advent of the “synthetic data age”, our work signals the end of current neural scaling laws and opens the door to a puzzling array of new phenomena in a world where the training corpora are enriched with AI generated data. We demonstrate that scaling laws cease to persist; test error tapers off due to altered, less heavy tailed, data distributions. Yet, new opportunities arise from careful mixture and data curation, as we have shown, with interesting effects at the interplay of clean and synthesized data. We must recognize new learning plateaus and, for instance, adjust to changed learning curves from blending clean and synthetic data to unintended early stopping. A notable feature of our work is that our theory is effective - we observe the predicted phenomena for relevant large models in two different settings.
Our contributions call for a more responsible, or “collapse-aware”, proliferation of synthesized data. Scale is not all you need: more work on watermarking for synthetic data is needed, to make it more distinguishable from original human-annotated data. “Real” data will become an even more valuable resource in the future, as we are ushering in the “beyond scaling” era.
Acknowledgements
@@
Impact Statement
This study contributes to the discourse on AI scaling laws by tackling the emergent challenge presented by synthetic data within training datasets. It holds particular relevance for fields that depend heavily on data accuracy, such as natural language processing and generative AI model development. We investigate the concept of ”model collapse” as a cautionary tale, highlighting the potential risks to sustainable model performance in the face of an increasing reliance on synthetic data. Through a combination of theoretical analysis and empirical evidence, we advocate for the strategic integration of clean data to mitigate these risks and enhance the robustness of AI systems. Our findings carry significant implications for industries where decision-making is progressively data-driven, suggesting a need for a shift in data management strategies to emphasize the preservation of model integrity and longevity. Additionally, this research highlights a looming issue: the potential scarcity of clean data. It underscores the urgency of focusing on the cultivation and preservation of high-quality datasets for AI training, to safeguard the future of AI development.
@@ Appendix A Prior Work
LeBrun et al. (2021) first investigate training on AI-generated texts using transformers and LSTMs, revealing the distributional distortion inherent in these neural language models. The phenomenon of model collapse is first proposed by Shumailov et al. (2023) and has recently appeared in the literature in the context of language and image generation. Several recent works demonstrate facets of this phenomenon empirically in various settings (Hataya et al., 2023; Martínez et al., 2023a, b; Bohacek & Farid, 2023; Briesch et al., 2023; Guo et al., 2023; Fan et al., 2023). Only few recent works also provide some accompanying theoretical analysis (Shumailov et al., 2023; Alemohammad et al., 2023; Bertrand et al., 2023) which we outline now.
@@
Shumailov et al. (2023) define model collapse and attribute it to two mechanisms: finite sampling when training a model (leading to cut off of low-probability data) and function approximation errors (the model is not sufficiently expressive to model the true distribution). They observe (and, for a single Gaussian, prove) that upon iteratively resampling finite “training data” the generated distribution becomes more and more peaked. Other models studied empirically are mixtures of (two) Gaussians and VAEs on MNIST. To study language models, Shumailov et al. (2023) iteratively fine tune Meta’s OPT-125M model on wikidata2. For generation of new text they use a 5-way beam search, which, by its nature, (approximatively) generates only low-perplexity data.
@@
Alemohammad et al. (2023) conduct an empirical and analytical analysis on generative image models of what they term the “self-consuming” or “autophaguous” loop. They conclude that without enough fresh real data at each generation, future models necessarily will have their precision or recall decrease. They model the influence of each new AI-generation via a generic sampling bias $0\leq\lambda\leq 1$. In the case of image generation this refers to feature parameters at generation that favor quality over diversity (suitably quantified). More precisely, $\lambda=1$ corresponds to unbiased sampling and $\lambda=0$ corresponds to sampling from the modes of the generative distribution. $\lambda$ models biased sampling methods commonly used in generative modeling practice, such as truncation in BigGAN and StyleGAN or guidance in diffusion models. In the case of Gaussian distributions, $\lambda$ is the shrinking factor of the variance of the next generation. Their empirical work studies GANs and denoising diffusion probabilistic models for image generation on FFHQ and MNIST and single Gaussians for both theoretical and empirical observations. As in (Shumailov et al., 2023) they observe (and prove for the case of a single Gaussian) that estimation error alone leads to vanishing variance with number of iterations. Alemohammad et al. (2023) also empirically observe an initial boost in performance in a regime where modest amounts of synthetic data are mixed with the original data before larger amounts of synthetic data lead to ultimate degradation. This might mimick larger-scale results that demonstrate how synthetic data mixed with true data improves performance in some scenarios (see Benefits of synthesized data below). Indeed, in its simplest form, data augmentation (rotations, cropping etc. ), a widespread highly beneficial practice in ML training, can be viewed as the simplest form of data generation.
@@
Let us mention the study of Bertrand et al. (2023) in the context of image generation, which sheds light on the critical role of data composition in the stability and effectiveness in generative models. They explore scenarios involving a mix of clean data, representative of the true distribution, and synthesized data from previous iterations of the generator. Their analysis reveals that if the data mix consists exclusively of synthesized data, the generative process is likely to degenerate over time, leading to what they describe as a ‘clueless generator’. Thus, the generator collapses: it progressively loses its ability to capture the essence of the data distribution it was intended to model. Conversely, they found that when the proportion of clean data in the mix is sufficiently high, the generator, under certain technical conditions, retains the capability to learn and accurately reflect the true data distribution. This work sheds light on the critical role of data composition in the stability and effectiveness of generative models.
@@
Several empirical studies confirm the deleterious effect of training on self-generated data: In the context of image generation, Martínez et al. (2023a, b) report degradation of models trained on AI-generated data. Specifically, they use a Denoising Diffusion Implicit Model and a few (relatively small) datasets (e.g. Orchids, MNIST) to demonstrate visual degradation when training in successive generations of AI-generated data. Hataya et al. (2023) “conclude that generated images negatively affect downstream performance, while the significance depends on tasks and the amount of generated images”, Bohacek & Farid (2023) reports that the popular StableDiffusion model collapses when iteratively retrained on self-generated faces, even with as little as $3\%$ synthetic data mixed into the original training set. For text, Briesch et al. (2023) use nanoGPT [^5] https://github.com/karpathy/nanoGPT on a curated 10K logical-expression dataset to demonstrate the iterative collapse of self-consuming loops - the model and dataset are sufficiently small to allow training from scratch. Guo et al. (2023) observe a decline in linguistic diversity metrics across iteratively fine-tuned LLMs.
@@
To our knowledge, rigorous theory (or even empirical demonstrations) on mitigation strategies against model collapse are yet to come, with one notable exception in (Bertrand et al., 2023) (see below). Several works discuss the need for detection of AI-generated images or text (to avoid retraining on them), for example motivating research into watermarking strategies. Bertrand et al. (2023) analyze iterative retraining on a mixture of synthesized and original data under several technical assumptions and find that there are fixed points governing the stability of iterative retraining.
@@
There is a range of results showing benefits of AI-synthesized data in training better models, though mostly these results pertain to image data, specifically in the context of diffusion models (Azizi et al., 2023; He et al., 2023; Shipard et al., 2023; Bansal & Grover, 2023; Lin et al., 2023), though not only (see Dai et al. (2023); Xu et al. (2023); Huang et al. (2022); Wang et al. (2023) for chat-related examples). One might argue that they either throw model-collapse caution to the winds or, possibly, settle in the protected corner where mild amounts of synthetic data (or larger amounts of “mildly synthetic” data, like in the case of data augmentation) helps. In particular, often benefits of synthetic data are observed when the synthetic data is generated by a model trained for a different use case than the downstream task (like images synthesized from diffusion models helping classification models) or generated by a stronger model (He et al., 2023; Shipard et al., 2023; Bansal & Grover, 2023; Lin et al., 2023). However, other works critically analyze the purported benefit of generated data. Burg et al. (2023) find that while synthesized data from a diffusion model helps improving downstream tasks, such as classification, using the pre-training data of the diffusion model alone gives even stronger performance (which we can interpret as evidence of mild first-generation model collapse). All in all it is fair to say that the impact of data augmentation using generative models is still not fully understood.
@@
This study was extended to demonstrate many more power law relations in various scenarios (vision transformer, video modeling, multimodal models, and mathematical problem solving) (Henighan et al., 2021). In the machine translation (MT) setting, Gordon et al. (2021) quantify scaling laws for standard benchmarks like BLEU and explain them via cross-entropy power-law scaling, thus positing a first universality of scaling laws across metrics. Hernandez et al. (2021) demonstrate similar empirical power-law scaling for transfer learning and Aghajanyan et al. (2023) provide a vast experimental body of evidence for scaling laws in mixed-modal language models.
However, a few results have nuanced the view of scaling as a panacea to improved loss. For instance, McKenzie et al. (2023) present evidence for ”inverse sclaing” where flaws in the training objective or the data lead to U-shaped scaling.
@@
From a theoretical angle, scaling laws have been shown analytically even before the emergence of large foundation models. For instance, Caponnetto & de Vito (2007) characterize the power-law generalization error of regularized least-squares kernel algorithms. The role of optimization can also be taken into account in this setting (Nitanda & Suzuki, 2021). In the nonparametric literature, for example Schmidt-Hieber (2017) and Suzuki (2019) derived the test error scaling of deep neural network in fitting certain target functions and (Bordelon et al., 2020) analyze spectral dependence.
More recently, scaling laws have been shown for kernel models under the Gaussian design, e.g. in (Spigler et al., 2020; Cui et al., 2021, 2022) for regression and (Cui et al., 2023) for classification. Maloney et al. (2022) study scaling laws for the random feature model in the context of regression. In the context of memorization for heavy-tailed data scaling laws have been shown in the infinite-memory setting (Hutter, 2021), for ”quantized” skills (Michaud et al., 2023) and for certain random data-generation processes (Debowski, 2023). When taking model capacity and optimization into account, Cabannes et al. (2023) recently proved scaling laws in constraint-capacity associative memories.
@@
Scaling laws give us an insight on bang-for-the-buck style trade-off for model training. However, cross-entropy loss is not a goal in and of itself: we want to train models that are endowed with a larger and larger skill set as we scale them up. For instance, Gordon et al. (2021) provide intuition and empirics for the scaling of BLEU score for MT with cross-entropy loss as